If f(x)=2x^2-5x-3, then f(-3)=
A)0
B)30
C)18
D)48
Please show me step by step how this works

Answer by checkley71(8403)

F(X)=2X~2-5X-3 & F(-3) OR 2(-3)~2-5(-3)-3=2*18+15-3=33-3=30 (B)30